In the absence of air resistance, all objects fall at the same rate according to

                                   y = vi t +1 / 2 g t2

where vi is the initial velocity of the falling object. If dropped from rest, the equation becomes

y = 1 / 2 g t2

For a 2 m release point, calculate the impact velocity  

  1. if dropped from rest
  1. If released with a downward velocity of 2.00 m/s.

       C. if released with an upward velocity of 2.00 m/s.
Answer #1

Part A.

Let the impact velocity is v

In Uniform accelerated motion

v2=vi2+2gd

In this case vi=0 and d=2m

so v2=0+2(9.8)2

v=6.26 m/s

Part B.

if vi=2m/s

then

v2=22+2(9.8)2

v=6.57 m/s

Part c.

if vi=2 m/s in upward direction then at the releasing point it is 2m/s in downward diection

so this case is same as case2

v=6.57m/s